OpenStudy (anonymous):

which of the following is the correct notation for complex number 76+√49??

OpenStudy (saifoo.khan):

\[\sqrt {49} \to \sqrt{7 \times 7} \to 7\] THerefore, 76 + sqrt 49 76 + 7 Solve.

OpenStudy (anonymous):

sorry i forgot to mention a "-" beside the sqrt of 49 ... o_o

OpenStudy (unklerhaukus):

inside or outside?

OpenStudy (anonymous):

inside

OpenStudy (anonymous):

√-49

OpenStudy (unklerhaukus):

\[76+\sqrt{-49}\]\[=76+\sqrt{-1\times49}\]\[=76+\sqrt{-1}\times\sqrt{49}\]=76+7i\]